What is the LHV of n-pentane at standard condition in KJ/kgmole n-pentane? (n-pentane --gas)
Answer #1

The balanced combustion reaction of n-pentane

C5H12 (g) + 8O2(g) = 5CO2 (g) + 6H2O (l)

Enthalpy change for the reaction = sum of Enthalpy of formation of products - sum of Enthalpy of formation of reactants

Hc = 5*Hf(CO2) + 6*Hf(H2O) - 8*Hf(O2) - Hf(C5H12)

= 5*(-393520)+ 6*(-285820) - 8*0 - (-146760)

= - 1967600 - 1714920 + 146760

= - 3535760 kJ/kmol

Higher heating value of gaseous n-pentane = - (-3535760)

HHV = 3535760 kJ/kmol

LHV = HHV - moles of H2O * enthalpy of vaporization of H2O

= (3535760 kJ/kmol) - (6 kmol H2O / kmol C5H12) * (44010 kJ/kmol H2O)

= 3271700 kJ/kmol n-pentane